# Modeling Sustainable City Trips: Integrating CO2e Emissions, Popularity, and Seasonality into Tourism Recommender Systems

### v2

@@ 1. Introduction
Recommender Systems (RS) provide tailored content to individual preferences, spanning diverse domains like e-commerce, social media, news, and more, effectively managing information to prevent overload (Abdollahpouri et al., 2020). In travel and tourism, RS is pivotal in simplifying trip planning by providing personalized recommendations for destinations, accommodations, activities, and more (Isinkaye et al., 2015). However, this is a particularly challenging domain due to the influence of dynamic factors such as seasonality and travel regulations (Balakrishnan and Wörndl, 2021), as well as constraints related to capacity-limited resources such as airline seats, hotel rooms, and event tickets (Abdollahpouri and Burke, 2021).
@@
Traditionally, RS focused on delivering accurate user recommendations, but in practice, they function as a convergence point for multiple stakeholders, making it a multistakeholder scenario (Abdollahpouri et al., 2020). Recognizing the interests of all stakeholders becomes crucial in this dynamic. Our stakeholder classification, inspired by Balakrishnan and Wörndl (2021), identifies four key categories: consumers, item providers, platform, and society, aligning with common touristic recommendation scenarios. Despite this seemingly straightforward categorization, real-world stakeholder relationships are often more intricate. Each stakeholder is vested in the traveler’s journey, and optimizing consumer recommendations can yield benefits for all involved parties (Abdollahpouri et al., 2020). Complexities arise when the goals of stakeholders conflict, mainly as profit motives often drive them, leading to inevitable trade-offs in achieving fairness in these systems (Jannach and Bauer, 2020). This necessitates adopting a multistakeholder approach, acknowledging stakeholders’ interdependence, and balancing their objectives when designing fair Tourism Recommender Systems (TRS).
@@
Tourism’s impact goes beyond active participants; it affects the local environment and businesses and profoundly influences the balance of nature. Thus, developing a fair TRS involves recommending sustainable options and fostering responsible tourism practices. World Tourism Organization and United Nations Development Programme define sustainable tourism as ”tourism that takes full account of its current and future economic, social and environmental impacts, addressing the needs of visitors, the industry, the environment, and host communities”  (Gössling, 2017). Achieving sustainability in tourism requires interventions at various levels, including municipal policies and regulations (Werthner et al., 2015). However, measuring sustainability at destinations poses a significant challenge, impeding effective decision-making, management, and meeting destination needs (Fernández and Rivero, 2009). In the context of tourism, a destination is characterized by ”a country, state, region, city, or town that is actively marketed or markets itself as an appealing place for tourists to visit” (Beirman, 2020). A destination’s sustainability is crucial for long-term competitiveness and visitor satisfaction, and should not solely be determined by arrival numbers or bed nights (Önder et al., 2017).
@@
One of the essential interventions where a well-designed TRS can play a vital role is regulating the number of tourists. A TRS can be especially helpful in addressing the challenges of over- and undertourism, both of which are on the rise due to factors such as low-cost aviation, affordable transportation, social media influence, and platforms like Airbnb  [^1] https://www.airbnb.com/ (Gowreesunkar and Vo Thanh, 2020). Overtourism, witnessed in popular destinations like Venice, Barcelona, Rome, and Dubrovnik, poses threats to historic preservation, the environment, residents, and overall tourist experiences, making it challenging to find reasonably priced housing in these cities (Dastgerdi and De Luca, 2023; Dodds and Butler, 2019). Conversely, undertourism, prevalent in under-explored destinations, results from a lack of infrastructure, publicity, and accessibility. Both scenarios have adverse consequences. For instance, the recent COVID-19 pandemic highlighted the adverse effects of undertourism, causing significant disruptions to the tourism and hotel industries (Galí, 2022). To address these issues, a TRS must be designed to provide responsible recommendations, considering the interests of all stakeholders. These systems should advocate for sustainable tourism practices, promoting responsible tourism while offering personalized suggestions to users. This involves recommendations encouraging tourists to visit destinations with minimal environmental impact, promoting less popular yet attractive locations, and balancing the tourist load uniformly throughout the year.
@@
A substantial amount of research has been conducted on developing fair recommendation systems that consider the interests of all stakeholders involved in tourism  (Rahmani et al., 2022; Shen et al., 2021; Weydemann et al., 2019; Wu et al., 2021). However, there has been limited focus on generating sustainable recommendations (Banerjee et al., 2023). This paper explores the concept of modeling Societal Fairness or S-Fairness. It emphasizes the impact of tourism on individuals who are not directly involved, such as residents, environment, and other stakeholders, collectively referred to as ”society” (Banerjee, 2023). These stakeholders often encounter challenges such as rising housing prices, environmental pollution, and traffic congestion due to heightened tourism activities in their vicinity.
@@ 2. Related Work
In the context of this study, we explore the related work along two dimensions — City Trip Recommendations in Section 2.1 and Sustainability in Tourism Recommender Systems in Section 2.2. This reflects the shift in recommender systems towards integrating personalization and user preferences in city trip recommendations while considering sustainability and societal fairness.
2.1. City Trip Recommendations
@@
City trip recommender systems are essential not only to simplify trip planning for users but also to cater to the evolving needs and challenges in the travel domain. Traditionally, recommender systems have predominantly centered on modeling user preferences to offer personalized suggestions. However, this domain presents distinctive challenges, including the intangibility of recommended items and dynamic factors such as seasonality, travel regulations, and resource constraints (Balakrishnan and Wörndl, 2021; Werthner and Ricci, 2004). Recent research has emphasized addressing these aspects as well.
@@
Besides techniques like collaborative filtering leveraging users’ past activities, similarities with other users, and network-based preferences are also prevalent in literature (Lu et al., 2012; Dadoun et al., 2019; Pu et al., 2020). Constraints defined by users, such as budget or time preferences, have also been addressed as critical components in tailoring travel packages (Xie et al., 2010; Lim et al., 2018). However, challenges like the cold start problem and data sparsity have led to the adoption of content-based approaches. These approaches construct domain models using relevant features for tourism, often derived from expert opinions, literature, or data-driven methods (Liu et al., 2011; Pu et al., 2020).
@@
While expert-driven models offer nuanced insights, their cost and complexity necessitate complementing them with diverse data sources, such as Location-based Social Networks (LBSNs) for venues (Lu et al., 2012; Dadoun et al., 2019). For instance, the data-driven system introduced by Roy and Dietz (2021) tailors composite city trips to users based on their preferences and constraints using data from the LSBNs. It suggests personalized itineraries for users by gathering information such as their home region, travel duration, and venue preferences. Similarly, the work by (Myftija and Dietz, 2020) uses a data-driven characterization of cities inside a conversational recommender system to recommend cities as destinations.  (Dietz and Weimert, 2018) present wOndary, a platform designed for global trip planning and sharing via crowdsourcing. Using content-based recommendation methods and a structured itinerary representation, wOndary tackles issues surrounding item discovery and routing in tourist trip design.  Massimo and Ricci (2019), explores the concept of clustering users with similar POI visit trajectories and then constructing a general user behavior model via inverse reinforcement learning. This model allows for generating recommendations based on learned behavioral patterns.
2.2. Sustainability in Tourism Recommender Systems
@@
Measuring Sustainability: Sustainable tourism presents complexities in defining objectives and indicators for a sustainable TRS. Ko (2005) assert that sustainability concerns vary by destination, necessitating specific dimensions and methodologies tailored to each location. However, this tailored approach raises comparability issues, as different sets of indicators hinder meaningful comparisons between destinations. In response,  Cernat and Gourdon (2012) develop the Sustainable Tourism Benchmarking Tool (STBT), comprising 54 indicators across seven dimensions. Though tested in 75 countries, the tool performs well only in Indonesia, Malaysia, and Thailand, where data availability is higher. Önder et al. (2017) suggest prioritizing the analysis of existing sustainable tourism indicators over introducing new, less practically applicable measures. They propose utilizing data envelopment analysis (DEA) on existing tourism information systems to model destination competitiveness and create a utility function sustainability index. DEA enables the identification of a destination’s efficiency and facilitates benchmarking against others to pinpoint areas for improvement. While this index is valuable for guiding policymakers and city developers in enhancing specific aspects of cities, it has not been integrated into TRS yet.
@@
Gorantla and Bansal (2023) utilize the Circles of Sustainability framework to assess a city’s sustainability across four domains and seven subdomains each, aggregating data sources to compute a simplified sustainability index, providing valuable insights into city sustainability in comparison to others. In contrast, Hoffmann et al. (2022) adopt a fully data-driven approach, analyzing Tripadvisor data to compare sustainability measures among hotels. Their unsupervised statistical learning approach yields improved performance in classifying sustainable hotels. However, the model’s explainability is limited; while it indicates larger hotels are more likely to be sustainable, causality cannot be explicitly inferred.
@@
Sustainable Recommendations: While numerous studies explore the quantification of sustainability in tourism, there is a notable scarcity of their application within recommender systems in this domain. Sustainable recommender system seeks to create a balance across economic, environmental, and social dimensions which is essential for creating a resilient ecosystem (Gorantla and Bansal, 2023). One can initiate the integration of features that capture user interest while simultaneously promoting sustainability. However, a sustainable recommender system should not be fully user-centric. The tourism industry involves different stakeholders, including consumers, providers, platforms, and society, thus it is natural for sustainable RS to implement a multistakeholder approach. In a recent examination of multistakeholder fairness within tourism recommender systems, it became apparent that existing studies predominantly concentrate on provider and consumer stakeholders. Surprisingly, the societal aspect, despite bearing significant impacts from tourism, is often overlooked as a stakeholder (Banerjee et al., 2023).
@@
A recent study by Merinov et al. (2022) introduces a multistakeholder model that not only considers user preferences but also the occupancy level of the destination. In another study by Patro et al. (2020), the multi-objective model focuses on providers’ sustainability by maintaining their exposure while also preventing overcrowding. Pachot et al. (2021) added local authorities as stakeholders, with economic growth, productive resilience, prioritizing basic necessities, and greener production as its objective. While Banik et al. (2023) address overcrowding by promoting fair recommendations that encourage users to choose greener alternatives, they lack a specific method for measuring the generation of recommended items.
@@
3. Destinations, Transportation, and Emission Estimations
Tourism contributes to approximately 8% of global CO2e emissions, stemming from various sources such as accommodation, food consumption, shopping, services, and agriculture. However, research has shown that transportation modes to and from destinations are the most significant contributors to the CO2e footprint, making up 49% of emissions. This highlights the importance of considering the emissions from transportation as one of the key sustainability indicators in tourism, especially for city policymakers (Lenzen et al., 2018). These transportation-related emissions often result from a combination of modes chosen by travelers, the distance traveled, and the average length of stay (Önder et al., 2017). For example, longer-distance trips that necessitate air travel between cities tend to incur higher emissions, posing a more significant environmental impact than shorter trips feasible with public transport like trains. Conversely, a short-distance trip by car may result in more emissions than a train.
@@ 3. Destinations, Transportation, and Emission Estimations
3.1. Extracting Destinations
In our scenario, a traveler is seeking a suitable city to visit for vacation from a specified starting location. They are presented with a list of European cities as potential destinations. The initial step involves collecting data on potential destinations for city trips. Our item space comprises 200 European cities or destinations spanning 43 countries. We have chosen these destinations due to the continent’s extensive connectivity via various modes of transportation, including flights, rail, and road, making it a highly popular destination among tourists. Data for European cities is sourced from the world cities database (World Cities Database, 2023), filtering for the top 200 most populated cities, each featuring at least one airport. Subsequently, we calculate transportation emissions for travel to each city using three modes — flights, driving, and rail where applicable, as explained below. Figure 1 illustrates the geographical dispersion of the 200 European cities considered for our analysis. It also shows the subset of cities specifically chosen for in-depth examination regarding driving and train connections in this study.
Figure 1. The geographical distribution of 200 European cities with at least one airport is depicted in blue, cities prioritized for driving are highlighted in orange, and those considered for their train networks are shown with a green +
3.2. Data Gathering: Transportation
@@ 3.2.1. Flights
We classified each route based on the Eurocontrol’s  [^5] https://www.eurocontrol.int/ definition for distances, segmenting them into four categories: very short haul, short haul, medium haul, and long haul. Specifically, very short-haul flights covered less than 500 km, short-haul flights ranged from 500 to 1,500 km, medium-haul flights spanned 1,500 to 4,000 km, and long-haul flights exceeded 4,000 km in flying distance. The scraped data from Google Flights lacked information on the distances between the two cities. Therefore, we employed the Great Circle Distance (GCD) (Wikipedia, 2023a, b) measurement to calculate the distances for each route. As indicated in Table 1, our dataset predominantly comprised short-haul and medium-haul flights, with a comparatively smaller number of long-haul flights.
3.2.2. Driving
@@
3.2.3. Trains
The railway network across Europe exhibits a diverse and country-specific management infrastructure. Unfortunately, no open-source API or standardized pan-European platform would enable us to aggregate data seamlessly from various railway networks. Consequently, our approach relied on country-specific rail networks, focusing on Germany due to easier data accessibility. We utilized web scraping techniques on the Bahn.Expert website to gather the necessary information. This platform uses the Deutsche Bahn APIs  [^9] https://data.deutschebahn.com/dataset.groups.apis.html to provide valuable insights into past and present data on trains and stations, particularly for Deutsche Bahn (DB) or German trains (Clasen, 2020).

### v3

@@ 1 Introduction
Recommender Systems (RS) provide tailored content to individual preferences, spanning diverse domains like e-commerce, social media, news, and more, effectively managing information to prevent overload [2]. In travel and tourism, RS is pivotal in simplifying trip planning by providing personalized recommendations for destinations, accommodations, activities, and more [40]. However, this is a particularly challenging domain due to the influence of dynamic factors such as seasonality and travel regulations [5], as well as constraints related to capacity-limited resources such as airline seats, hotel rooms, and event tickets [1].
@@
Traditionally, RS focused on delivering accurate user recommendations, but in practice, they function as a convergence point for multiple stakeholders, making it a multistakeholder scenario [2]. Recognizing the interests of all stakeholders becomes crucial in this dynamic. Our stakeholder classification, inspired by Balakrishnan and Wörndl [5], identifies four key categories: consumers, item providers, platform, and society, aligning with common touristic recommendation scenarios. Despite this seemingly straightforward categorization, real-world stakeholder relationships are often more intricate. Each stakeholder is vested in the traveler’s journey, and optimizing consumer recommendations can yield benefits for all involved parties [2]. Complexities arise when the goals of stakeholders conflict, mainly as profit motives often drive them, leading to inevitable trade-offs in achieving fairness in these systems [42]. This necessitates adopting a multistakeholder approach, acknowledging stakeholders’ interdependence, and balancing their objectives when designing fair Tourism Recommender Systems (TRS).
@@
Tourism’s impact goes beyond active participants; it affects the local environment and businesses and profoundly influences the balance of nature. Thus, developing a fair TRS involves recommending sustainable options and fostering responsible tourism practices. World Tourism Organization and United Nations Development Programme define sustainable tourism as "tourism that takes full account of its current and future economic, social and environmental impacts, addressing the needs of visitors, the industry, the environment, and host communities"  [33]. Achieving sustainability in tourism requires interventions at various levels, including municipal policies and regulations [77]. However, measuring sustainability at destinations poses a significant challenge, impeding effective decision-making, management, and meeting destination needs [24]. In the context of tourism, a destination is characterized by "a country, state, region, city, or town that is actively marketed or markets itself as an appealing place for tourists to visit" [9]. A destination’s sustainability is crucial for long-term competitiveness and visitor satisfaction, and should not solely be determined by arrival numbers or bed nights [54].
@@
One of the essential interventions where a well-designed TRS can play a vital role is regulating the number of tourists. A TRS can be especially helpful in addressing the challenges of over- and undertourism, both of which are on the rise due to factors such as low-cost aviation, affordable transportation, social media influence, and platforms like Airbnb  [^1] https://www.airbnb.com/ [34]. Overtourism, witnessed in popular destinations like Venice, Barcelona, Rome, and Dubrovnik, poses threats to historic preservation, the environment, residents, and overall tourist experiences, making it challenging to find reasonably priced housing in these cities [17, 20]. Conversely, undertourism, prevalent in under-explored destinations, results from a lack of infrastructure, publicity, and accessibility. Both scenarios have adverse consequences. For instance, the recent COVID-19 pandemic highlighted the adverse effects of undertourism, causing significant disruptions to the tourism and hotel industries [26]. To address these issues, a TRS must be designed to provide responsible recommendations, considering the interests of all stakeholders. These systems should advocate for sustainable tourism practices, promoting responsible tourism while offering personalized suggestions to users. This involves recommendations encouraging tourists to visit destinations with minimal environmental impact, promoting less popular yet attractive locations, and balancing the tourist load uniformly throughout the year.
@@
A substantial amount of research has been conducted on developing fair recommendation systems that consider the interests of all stakeholders involved in tourism  [64, 70, 78, 84]. However, there has been limited focus on generating sustainable recommendations [7]. This paper explores the concept of modeling Societal Fairness or S-Fairness. It emphasizes the impact of tourism on individuals who are not directly involved, such as residents, environment, and other stakeholders, collectively referred to as "society" [6]. These stakeholders often encounter challenges such as rising housing prices, environmental pollution, and traffic congestion due to heightened tourism activities in their vicinity.
@@ 2 Related Work
In the context of this study, we explore the related work along two dimensions – City Trip Recommendations in Section 2.1 and Sustainability in Tourism Recommender Systems in Section 2.2. This reflects the shift in recommender systems towards integrating personalization and user preferences in city trip recommendations while considering sustainability and societal fairness.
2.1 City Trip Recommendations
@@
City trip recommender systems are essential not only to simplify trip planning for users but also to cater to the evolving needs and challenges in the travel domain. Traditionally, recommender systems have predominantly centered on modeling user preferences to offer personalized suggestions. However, this domain presents distinctive challenges, including the intangibility of recommended items and dynamic factors such as seasonality, travel regulations, and resource constraints [5, 76]. Recent research has emphasized addressing these aspects as well.
@@
Besides techniques like collaborative filtering leveraging users’ past activities, similarities with other users, and network-based preferences are also prevalent in literature [50, 16, 63]. Constraints defined by users, such as budget or time preferences, have also been addressed as critical components in tailoring travel packages [85, 48]. However, challenges like the cold start problem and data sparsity have led to the adoption of content-based approaches. These approaches construct domain models using relevant features for tourism, often derived from expert opinions, literature, or data-driven methods [49, 63].
@@
While expert-driven models offer nuanced insights, their cost and complexity necessitate complementing them with diverse data sources, such as Location-based Social Networks (LBSNs) for venues [50, 16]. For instance, the data-driven system introduced by Roy and Dietz [66] tailors composite city trips to users based on their preferences and constraints using data from the LSBNs. It suggests personalized itineraries for users by gathering information such as their home region, travel duration, and venue preferences. Similarly, the work by [53] uses a data-driven characterization of cities inside a conversational recommender system to recommend cities as destinations.  [19] present wOndary, a platform designed for global trip planning and sharing via crowdsourcing. Using content-based recommendation methods and a structured itinerary representation, wOndary tackles issues surrounding item discovery and routing in tourist trip design.  Massimo and Ricci [51], explores the concept of clustering users with similar POI visit trajectories and then constructing a general user behavior model via inverse reinforcement learning. This model allows for generating recommendations based on learned behavioral patterns.
2.2 Sustainability in Tourism Recommender Systems
@@
Measuring Sustainability: Sustainable tourism presents complexities in defining objectives and indicators for a sustainable TRS. Ko [44] assert that sustainability concerns vary by destination, necessitating specific dimensions and methodologies tailored to each location. However, this tailored approach raises comparability issues, as different sets of indicators hinder meaningful comparisons between destinations. In response,  Cernat and Gourdon [13] develop the Sustainable Tourism Benchmarking Tool (STBT), comprising 54 indicators across seven dimensions. Though tested in 75 countries, the tool performs well only in Indonesia, Malaysia, and Thailand, where data availability is higher. Önder et al [54] suggest prioritizing the analysis of existing sustainable tourism indicators over introducing new, less practically applicable measures. They propose utilizing data envelopment analysis (DEA) on existing tourism information systems to model destination competitiveness and create a utility function sustainability index. DEA enables the identification of a destination’s efficiency and facilitates benchmarking against others to pinpoint areas for improvement. While this index is valuable for guiding policymakers and city developers in enhancing specific aspects of cities, it has not been integrated into TRS yet.
@@
Gorantla and Bansal [32] utilize the Circles of Sustainability framework to assess a city’s sustainability across four domains and seven subdomains each, aggregating data sources to compute a simplified sustainability index, providing valuable insights into city sustainability in comparison to others. In contrast, Hoffmann et al [38] adopt a fully data-driven approach, analyzing Tripadvisor data to compare sustainability measures among hotels. Their unsupervised statistical learning approach yields improved performance in classifying sustainable hotels. However, the model’s explainability is limited; while it indicates larger hotels are more likely to be sustainable, causality cannot be explicitly inferred.
@@
Sustainable Recommendations: While numerous studies explore the quantification of sustainability in tourism, there is a notable scarcity of their application within recommender systems in this domain. Sustainable recommender system seeks to create a balance across economic, environmental, and social dimensions which is essential for creating a resilient ecosystem [32]. One can initiate the integration of features that capture user interest while simultaneously promoting sustainability. However, a sustainable recommender system should not be fully user-centric. The tourism industry involves different stakeholders, including consumers, providers, platforms, and society, thus it is natural for sustainable RS to implement a multistakeholder approach. In a recent examination of multistakeholder fairness within tourism recommender systems, it became apparent that existing studies predominantly concentrate on provider and consumer stakeholders. Surprisingly, the societal aspect, despite bearing significant impacts from tourism, is often overlooked as a stakeholder [7].
@@
A recent study by Merinov et al [52] introduces a multistakeholder model that not only considers user preferences but also the occupancy level of the destination. In another study by Patro et al [60], the multi-objective model focuses on providers’ sustainability by maintaining their exposure while also preventing overcrowding. Pachot et al [57] added local authorities as stakeholders, with economic growth, productive resilience, prioritizing basic necessities, and greener production as its objective. While Banik et al [8] address overcrowding by promoting fair recommendations that encourage users to choose greener alternatives, they lack a specific method for measuring the generation of recommended items.
@@
3 Destinations, Transportation, and Emission Estimations
Tourism contributes to approximately 8% of global CO2e emissions, stemming from various sources such as accommodation, food consumption, shopping, services, and agriculture. However, research has shown that transportation modes to and from destinations are the most significant contributors to the CO2e footprint, making up 49% of emissions. This highlights the importance of considering the emissions from transportation as one of the key sustainability indicators in tourism, especially for city policymakers [47]. These transportation-related emissions often result from a combination of modes chosen by travelers, the distance traveled, and the average length of stay [54]. For example, longer-distance trips that necessitate air travel between cities tend to incur higher emissions, posing a more significant environmental impact than shorter trips feasible with public transport like trains. Conversely, a short-distance trip by car may result in more emissions than a train.
@@ 3 Destinations, Transportation, and Emission Estimations
3.1 Extracting Destinations
In our scenario, a traveler is seeking a suitable city to visit for vacation from a specified starting location. They are presented with a list of European cities as potential destinations. The initial step involves collecting data on potential destinations for city trips. Our item space comprises 200 European cities or destinations spanning 43 countries. We have chosen these destinations due to the continent’s extensive connectivity via various modes of transportation, including flights, rail, and road, making it a highly popular destination among tourists. Data for European cities is sourced from the world cities database [83], filtering for the top 200 most populated cities, each featuring at least one airport. Subsequently, we calculate transportation emissions for travel to each city using three modes — flights, driving, and rail where applicable, as explained below. Figure 1 illustrates the geographical dispersion of the 200 European cities considered for our analysis. It also shows the subset of cities specifically chosen for in-depth examination regarding driving and train connections in this study.
Figure 1: The geographical distribution of 200 European cities with at least one airport is depicted in blue, cities prioritized for driving are highlighted in orange, and those considered for their train networks are shown with a green +
3.2 Data Gathering: Transportation
@@ 3.2.1 Flights
We classified each route based on the Eurocontrol’s  [^5] https://www.eurocontrol.int/ definition for distances, segmenting them into four categories: very short haul, short haul, medium haul, and long haul. Specifically, very short-haul flights covered less than 500 km, short-haul flights ranged from 500 to 1,500 km, medium-haul flights spanned 1,500 to 4,000 km, and long-haul flights exceeded 4,000 km in flying distance. The scraped data from Google Flights lacked information on the distances between the two cities. Therefore, we employed the Great Circle Distance (GCD) [79, 80] measurement to calculate the distances for each route. As indicated in Table 1, our dataset predominantly comprised short-haul and medium-haul flights, with a comparatively smaller number of long-haul flights.
3.2.2 Driving
@@
3.2.3 Trains
The railway network across Europe exhibits a diverse and country-specific management infrastructure. Unfortunately, no open-source API or standardized pan-European platform would enable us to aggregate data seamlessly from various railway networks. Consequently, our approach relied on country-specific rail networks, focusing on Germany due to easier data accessibility. We utilized web scraping techniques on the Bahn.Expert website to gather the necessary information. This platform uses the Deutsche Bahn APIs  [^9] https://data.deutschebahn.com/dataset.groups.apis.html to provide valuable insights into past and present data on trains and stations, particularly for Deutsche Bahn (DB) or German trains [14].